What is the slope of the line given below ?
y = 2/17x
Can You please Show work to How you got The answer Thank You

Answer by jim_thompson5910(35256)

The equation is in the form y = mx+b where m = 2/17 and b = 0


Since 'm' is the slope, the slope is 2/17.